A roller coaster ride in a park includes a vertical drop of 60m. Assume that the roller coaster has a speed of nearly zero when it reaches the top of the hill. Neglected friction, find the speed of the riders at the bottom of the hill. Show complete solution.


1
Expert's answer
2022-01-10T09:16:28-0500

The law of conservation of energy says

"E_i=E_f"

Total mechanical energy is a sum of kinetic and potential energy. So


"E_i=\frac{mv^2}{2}+0"

"E_f=0+mgh"

Therefore

"\frac{mv^2}{2}=mgh"

Finally

"v=\sqrt{2gh}=\sqrt{2*9.8*60}=34.5\:\rm m/s"